One day Tammy asked her mother on her birthday, "mom how old are you?" Her mother replied I am 12 years older than your aunt tamika and when I married your dad 20 years ago, I was twice as old as Aunt tamika?
can you solve problem in equation and tell me how you got to that answer

Answer by richwmiller(17219)

m=t+12;
m-20=2*(t-20)
m=44
t=32